Find the point of intersection of the pair of straight lines
8x + 8y=32
−9x + 7y= 4

Answer by shweta(56)

8x + 8y=32 ..(1)
-9x + 7y=4 ..(2)
Multiply equation 1 by 7 and equation 2 with 8
8x + 8y=32 *7
-9x + 7y=4 *8
56x + 56y= 224
-72x + 56y=32
Subtract both the equations
56x-(-72x) +56y -56y= 224-32
56x+ 72x= 192
128x= 192
x= 192/128
x= 3/2
Substitute the value of x in equation 1
8x+ 8y=32
8*(3/2) + 8y= 32
12 + 8y= 32
8y= 32-12
8y= 20
y= 20/8
y= 5/2
Point of intersection of these two lines is ( 3/2,5/2)